# Connecting Neural Models Latent Geometries with Relative Geodesic Representations

### v1

@@ Abstract
Neural models learn representations of high-dimensional data on low-dimensional manifolds. Multiple factors, including stochasticities in the training process, model architectures, and additional inductive biases, may induce different representations, even when learning the same task on the same data. However, it has recently been shown that when a latent structure is shared between distinct latent spaces, relative distances between representations can be preserved, up to distortions. Building on this idea, we demonstrate that exploiting the differential-geometric structure of latent spaces of neural models, it is possible to capture precisely the transformations between representational spaces trained on similar data distributions. Specifically, we assume that distinct neural models parametrize approximately the same underlying manifold, and introduce a representation based on the pullback metric that captures the intrinsic structure of the latent space, while scaling efficiently to large models. We validate experimentally our method on model stitching and retrieval tasks, covering autoencoders and vision foundation discriminative models, across diverse architectures, datasets, and pretraining schemes.
@@ 1 Introduction
Figure 1: Neural models trained on similar data learn parametrizations of the same manifold. NNs learn parametrizations ($D_{1},D_{2}$) of the same underlying manifold $\mathcal{Y}$ up to isometric transformations $\mathcal{T}$. By pulling back the metric from $\mathcal{Y}$, relative geodesic representations are invariant to the transformation $\mathcal{T}$ between latent spaces $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$
Neural models learn meaningful representations of high-dimensional data generalizing to many tasks, spanning different data modalities and domains. Recent research reveals that these models often develop similar internal representations when exposed to similar inputs Li et al. (2015); Moschella et al. (2023); Fumero et al. (2024); Kornblith et al. (2019a), a phenomenon that was observed in biological networks Laakso and Cottrell (2000); Haxby et al. (2001). Remarkably, even when models have different architectures, their internal representations can frequently be aligned through a simple, e.g., orthogonal, transformation Maiorca et al. (2024); Lähner and Moeller (2024a); Moayeri et al. (2023). This suggests a certain consistency in how neural nets encode information, emphasizing the importance of studying these internal representations, and the transformations that relate them, to the extent to hypothesize whether neural nets are converging toward a single representation of reality Huh et al. (2024).
One strategy to understand how different models are related is to identify representations that are invariant to transformations between distinct models’ representational spaces. A simple and effective recipe is that of relative representations Moschella et al. (2023), where samples are represented as a function of a fixed set of latent representations. The similarity function employed is cosine similarity, hinting at the fact that representations across distinct models are subject to angle preserving transformations. However, the choice of similarity function should not be limited to only capturing invariances of one class of transformations. As shown in Cannistraci et al. (2024); Fumero et al. (2021), other choices can be good as well, and there’s not a clear best choice among different transformations for capturing transformation across distinct latent spaces.We posit that when it is possible to relate distinct neural models’ representational spaces, this suggests that neural models are learning distinct parametrizations of the same underlying manifold (see Figure 1). In this paper, we employ geodesic distance in the latent space as a metric for relative representations. This approach ensures that the relative space remains approximately invariant to the isometries and reparametrization of the data’s manifold, as characterized by a Riemannian structure. Our contributions can be summarized as follows:
@@
  • We observe that different pullback metrics are suitable for different tasks, showing for the first time how to get meaningful pullback metrics from discriminative models, such as classifiers and self-supervised models.
  • We test relative geodesics on retrieval and stitching tasks on autoencoders and real vision foundation models, across different seeds, architectures, and training strategies, outperforming previous methods.
@@ 2 Related Work
Representation alignment.   Numerous studies have shown that neural networks trained under different initializations, architectures, or objectives learn highly similar internal feature representations Bonheme and Grzes (2022); Kornblith et al. (2019b); Klabunde et al. (2023); Li et al. (2015); Bengio et al. (2014); Maiorca et al. (2024); Huh et al. (2024); Guth et al. (2024); Chang et al. (2022); Conneau et al. (2018); Tsitsulin et al. (2020); Nejatbakhsh et al. (2024). This correspondence becomes stronger in wide and large networks Barannikov et al. (2022); Morcos et al. (2018); Somepalli et al. (2022). Leveraging these aligned embeddings, a simple linear transformation often suffices to map one network’s latent space onto another’s, enabling techniques such as model stitching, where components from different networks can be interchanged with minimal loss in performance Fumero et al. (2024); Bansal et al. (2021); Csiszárik et al. (2021). In practice, aligning two independently learned latent spaces often requires only a linear transformation, which achieves comparable downstream task performance (Moayeri et al., 2023; Merullo et al., 2023; Maiorca et al., 2024; Lähner and Moeller, 2024b).
Latent space geometry.   Early work on the geometry of deep latent representations focused on autoencoders, where the decoder’s mapping from latent to data space induces a natural pull-back metric under the assumption that the ambient space is Euclidean (Shao et al., 2018; Tosi et al., 2014; Arvanitidis et al., 2018). The Riemannian viewpoint allows one to compute geodesic paths and meaningful distances that respect the manifold structure of the learned embedding. Subsequent research has introduced computationally efficient approximations, such as energy-based proxies, and extended these ideas to estimate local curvature for improved interpolation and sampling (Chen et al., 2019; Chadebec and Allassonnière, 2022; Loaiza-Ganem et al., 2024; Arvanitidis et al., 2021, 2022a). In the context of discriminative models, one can obtain a Riemannian metric primarily using two approaches (Grosse, 2022), either by pulling back the Fisher Information Matrix (Amari, 2016; Arvanitidis et al., 2022b) or by assuming a Euclidean geometry on logit space and pulling back the metric.
@@ 3.1 Notation and Background
Neural networks (NNs) are parametric functions $F_{\theta}$, composed of an encoding map and a decoding map, represented as $F_{\theta}=D_{\theta_{2}}\circ E_{\theta_{1}}$. The encoder $E_{\theta_{1}}:\mathcal{X}\mapsto\mathcal{Z}$ generates a latent representation $z=E_{\theta_{1}}(x)$, where $x\in\mathcal{X}$ to the input domain $\mathcal{X}$, and the latent space $\mathcal{Z}$. The decoder $D_{\theta_{2}}$ is responsible for performing the task at hand, such as reconstruction or classification. For simplicity, we omit the parameter dependence ($\theta$) in our notation moving forward. For any single module $E$ (or equivalently $D$), we use $E_{\mathcal{X}}$ to denote that the module $E$ was trained on the domain $\mathcal{X}$. In the next sections, we will provide the necessary background to introduce our method.
@@
Latent Space Communication.   Given a pair of domains $\mathcal{X},\mathcal{Y}$, a pair of neural models trained on them $F_{\mathcal{X}}^{1},F_{\mathcal{Y}}^{2},$ and a partial correspondence between the domains $\Gamma:\mathcal{A}_{\mathcal{X}}\mapsto\mathcal{A}_{\mathcal{Y}}$ where $\mathcal{A}_{\mathcal{X}}\subset\mathcal{X}$ and $\mathcal{A}_{\mathcal{Y}}\subset\mathcal{Y}$, the problem of latent space communication is the one of finding a full correspondence $\Lambda:E^{1}(\mathcal{X})\mapsto E^{2}(\mathcal{Y})$ between the two domains, from $\Gamma$. In a simplified setting, e.g., two models trained with different initialization or architectures on the same data $\mathcal{X}=\mathcal{Y}$ and the correspondence is the identity. When $\mathcal{X}\neq\mathcal{Y}$, the problem becomes multimodal.
@@
Relative representations.   The relative representations framework Moschella et al. (2023) provides a straightforward approach to represent each sample in the latent space according to its similarity to a set of fixed training samples, denoted as anchors. Representing samples in the latent space as a function of the anchors corresponds to transitioning from an absolute coordinate frame into a relative one defined by the anchors and the similarity function. Given a domain $\mathcal{X}$, an encoding function $E_{\mathcal{X}}:\mathcal{X}\to\mathcal{Z}$, a set of anchors $\mathcal{A}_{\mathcal{X}}\subset\mathcal{X}$, and a similarity or distance function $d:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$, the relative representation for a sample $x\in\mathcal{X}$ is:
$RR(z;\mathcal{A}_{\mathcal{X}},d)=\bigoplus_{a_{i}\in\mathcal{A}_{\mathcal{X}}}d(z,E_{\mathcal{X}}(a_{i})),$
@@
where $z=E_{\mathcal{X}}(x)$, and $\bigoplus$ denotes row-wise concatenation. In the original method Moschella et al. (2023), $d$ was the cosine similarity. This choice induces a representation invariant to angle-preserving transformations. In this work, our focus is to leverage the intrinsic geometry of latent spaces to employ a metric that captures isometric transformations between data manifolds.
@@
Latent space geometry.   For the latent space of a neural network, it is generally hard to reason about its Riemannian structure. However, it is often easier to assign a Riemannian structure to the output space. As such, one can define a pullback metric from the output space to the latent space, which is a standard operation in Riemannian geometry (see ,e.g., Ch.2.4 of Do Carmo and Flaherty Francis (1992)).
@@
Formally, the decoder $D:\mathcal{Z}\mapsto\mathcal{X}$ takes as input a latent representation $z\in\mathcal{Z}$ and outputs $x$. Given a Riemannian metric defined on $x$ as $G_{\mathcal{X}}(x)$. Then, the Riemannian metric at $z$ is
$G_{\mathcal{Z}}(z)=\left(\frac{\partial x}{\partial z}\right)^{\!\!\top}G_{\mathcal{X}}(x)\left(\frac{\partial x}{\partial z}\right)=J_{z}(D)^{T}G_{\mathcal{X}}(x)J_{z}(D),$
@@
where $J_{z}(D)$ is the Jacobian of $D$ at $z$. The metric tensor $G_{\mathcal{X}}$ is useful to compute quantities such as lengths, angles, and areas on $\mathcal{M}$. Given a smooth curve $\gamma:[a,b]\mapsto\mathcal{M}$, its arc length is
$L(\gamma)$ $=\int^{b}_{a}\sqrt{v(t)^{\top}G(t)v(t)^{\top}}\mathrm{d}t.$ (1)
@@
A slight variation of the above functional gives the geodesic energy $\mathcal{E}$ of $\gamma$ (Arvanitidis et al., 2018; Shao et al., 2018)
$\mathcal{E}(\gamma)$ $=\frac{1}{2}\int^{b}_{a}v(t)^{\top}G(t)v(t)^{\top}\mathrm{d}t,$ (2)
@@
where $v(t)=\dot{\gamma}(t)$. Both can be discretized and approximated in practice using finite difference approaches (Yang et al., 2018; Shao et al., 2018). Geodesic paths minimize both the arc length and the energy, where the latter is usually preferred for numerical stability Hauberg (2025). The arc length instead has the property of being invariant to reparametrizations of the manifold,
Proposition 3.1 ((Do Carmo and Flaherty Francis, 1992)).
@@
Let $\gamma:[a,b]\to\mathcal{M}$ be a smooth curve on a Riemannian manifold $(\mathcal{M},G)$, and let $\varphi:[\alpha,\beta]\to[a,b]$ be any smooth, strictly increasing reparametrization. Define $\gamma(\tau)^{\prime}=\gamma\bigl{(}\varphi(\tau)\bigr{)}.$ Then the Riemannian length of $\gamma$ is unchanged:
[…]
@@ 3.2 Relative geodesics representations
1:Sample $x\in\mathcal{X}$, anchors $\mathcal{A}_{\mathcal{X}}$, encoder $E$, decoder $D$, metric $G_{\mathcal{X}}$, steps $N$, step size $\Delta t$, mode $\in\{\texttt{energy},\texttt{distance}\}$
2:$RR^{geo}(x;\mathcal{A}_{\mathcal{X}})$
3:$z\leftarrow E(x)$,  $RR^{geo}\leftarrow[\ ]$
@@
5:     $z_{a}\leftarrow E(a)$,  $d\leftarrow 0$
6:     for $j=1$ to $N$ do
7:         $\gamma_{j}\leftarrow(1-\tfrac{j}{N})z+\tfrac{j}{N}z_{a}$
8:         $\gamma_{j-1}\leftarrow(1-\tfrac{j-1}{N})z+\tfrac{j-1}{N}z_{a}$
9:         $v\leftarrow D(\gamma_{j})-D(\gamma_{j-1})$
10:         $G\leftarrow G_{\mathcal{X}}(D(\gamma_{j}))$
11:         $s\leftarrow v^{\top}Gv$
@@
When considering a differential geometry perspective, the problem of latent space communication can be interpreted as finding a transformation between the data manifolds $\mathcal{M}_{1},\mathcal{M}_{2}$ approximated by two neural models $F_{1},F_{2}$. The relative representation framework captures this transformation implicitly if equipped with the right metric. A natural candidate for this metric is the geodesic distance defined on $\mathcal{M}_{1},\mathcal{M}_{2}$, respectively. This choice makes the relative representations invariant to isometric transformation of the manifolds $\mathcal{M}_{1},\mathcal{M}_{2}$. However, for high-dimensional problems, the high cost of computing the geodesic renders the above methods inappropriate Shao et al. (2018); Chen et al. (2019). Furthermore, one can argue against directly using the latent geometry induced by deterministic models from a theoretical perspective (Hauberg, 2019), as it may result in undesirable properties, e.g., the resulting geodesics going outside of the data manifold.
@@
We therefore consider using the approximate energy of the straight line (in the Euclidean sense) connecting the representations in the latent space
$RR^{geo}(z;\mathcal{A}_{\mathcal{X}})=\bigoplus_{a_{i}\in\mathcal{A}_{\mathcal{X}}}\mathcal{E}(\tilde{\gamma}(z,E_{\mathcal{X}}(a_{i})))$
where $\tilde{\gamma}(z_{1},bz_{2})=(1-\alpha)z_{1}+\alpha z_{2}$ is the convex combination between the points $z_{1},z_{2}$.
It can be easily seen that for $\bar{\gamma}$ the three quantities are related by the following bounds:
$d(\mathbf{z}_{0},\mathbf{z}_{1})^{2}\leq L^{2}(\tilde{\gamma})\leq 2\mathcal{E}(\tilde{\gamma})$ (3)
On a Riemannian manifold, a natural choice to form such a representation is to use the Riemannian arc length of a curve defined respectively in Eq. 1 and the energy in Eq. 2.
Discretization.   The energy and arc length can be approximated using finite difference schemes,
$\mathcal{E}(\gamma)$ $=\sum_{i=1}^{N}E_{i}=\frac{1}{2}\sum_{i=1}^{N}v(t_{i})^{\top}G(t_{i})v(t_{i})\Delta t,$ (4)
$L(\gamma)$ $=\sum_{i=1}^{N}d_{i}=\sum_{i=1}^{N}\sqrt{v(t_{i})^{\top}G(t_{i})v(t_{i})}\Delta t,$ (5)
where $\Delta t=\frac{1}{N}$, with $N$ being the number of discretization steps.
When the step size is small enough, both quantities in the latent space can be approximated by their counterpart on the output space (Shao et al., 2018). For Euclidean geometry, the geodesic arc length is given in closed form as the geodesics are straight lines. Note that, unlike the energy, the curve length is invariant under reparametrizations (proposition 3.1). As such, we use the curve length.
@@
Approximate geodesic distances.   Our choice comes with three advantages (i) Efficiency: since we avoid gradient descent, the computation reduces to a single forward pass for each step in $\gamma$, (ii) The approach is minimal yet sufficient as we only need reasonably accurate estimates of the lengths rather than the entire geodesic trajectory, (iii) Since we don’t have to perform any optimizations we can use directly the arc length benefiting from its invariance to reparametrizations. To assess how close the straight line energy approximation (2) is to the true geodesic distances, we first encoded 100 samples (10 per class, sorted by label) from MNIST Deng (2012) and CIFAR-10 using a simple autoencoder’s encoder. We then computed pairwise distance matrices over these latent representations using both methods, and the results are displayed in Fig. 2. Visually, both distance matrices exhibit the same block-diagonal structure, mainly due to belonging to the same class, and clustering patterns. Numerically, their Spearman rank correlation exceeds 0.99 with only 8 discretization points (see Appendix for correlation results across different numbers of discretization steps and for implementation details).

### v2

@@ Abstract
Neural models learn representations of high-dimensional data on low-dimensional manifolds. Multiple factors, including stochasticities in the training process, model architectures, and additional inductive biases, may induce different representations, even when learning the same task on the same data. However, it has recently been shown that when a latent structure is shared between distinct latent spaces, relative distances between representations can be preserved, up to distortions. Building on this idea, we demonstrate that exploiting the differential-geometric structure of latent spaces of neural models, it is possible to capture precisely the transformations between representational spaces trained on similar data distributions. Specifically, we assume that distinct neural models parametrize approximately the same underlying manifold, and introduce a representation based on the pullback metric that captures the intrinsic structure of the latent space, while scaling efficiently to large models. We validate experimentally our method on model stitching and retrieval tasks, covering autoencoders and vision foundation discriminative models, across diverse architectures, datasets, pretraining schemes and modalities. Code is available at https://github.com/marc0git/RelativeGeodesics.
@@ 1 Introduction
Figure 1: Neural models trained on similar data learn parametrizations of the same manifold. NNs learn parametrizations ($D_{1},D_{2}$) of the same underlying manifold $\mathcal{Y}$ up to isometries $T$. Pulling back the metric from $\mathcal{Y}$ makes relative geodesic representations invariant to transformations $T$ between latent spaces $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$.
Neural models learn meaningful representations of high-dimensional data generalizing to many tasks, spanning different data modalities and domains. Recent research reveals that these models often develop similar internal representations given similar inputs (Li et al., 2015; Moschella et al., 2023; Fumero et al., 2024; Kornblith et al., 2019), a phenomenon that was observed in biological networks (Laakso and Cottrell, 2000; Haxby et al., 2001). Remarkably, even when models have different architectures, their internal representations can frequently be aligned through a simple, e.g. orthogonal, transformation (Maiorca et al., 2024; Lähner and Moeller, 2024; Moayeri et al., 2023). This suggests a certain consistency in how neural nets encode information, emphasizing the importance of studying the internal representations and the transformations that relate them, to the extent to hypothesize whether neural nets are converging toward a unique representation of reality (Huh et al., 2024).
One strategy to understand how different models are related is to identify representations that are invariant to transformations between distinct models’ representational spaces. A simple and effective recipe is that of relative representations (Moschella et al., 2023), where samples are represented as a function of a fixed set of latent representations. The similarity function employed is cosine similarity, hinting at the fact that representations across distinct models are subject to angle preserving transformations. However, the choice of similarity function should not be limited to only capturing invariances of one class of transformations. As shown in Cannistraci et al. (2024); Fumero et al. (2021), other choices can be good as well, and there is not a clear best choice among different transformations for capturing transformation across distinct latent spaces. We posit that when it is possible to relate distinct neural models’ representational spaces, neural models are learning distinct parametrizations of the same underlying manifold (see Figure 1). In this paper, we employ geodesic distance in the latent space for relative representations. This approach ensures that the relative space remains approximately invariant to the isometries and reparametrization of the data’s manifold, as characterized by a Riemannian structure. Our contributions can be summarized as follows:
@@
  • We show how to get meaningful pullback metrics from discriminative models, such as classifiers and self-supervised models.
  • We test relative geodesics on retrieval and stitching tasks on autoencoders and vision foundation models, across different models, training schemes, and modalities, outperforming prior methods.
@@ 2 Related Work
Representation alignment.   Numerous studies have shown that neural networks trained under different initializations, architectures, or objectives learn highly similar internal feature representations (Bonheme and Grzes, 2022; Kornblith et al., 2019; Klabunde et al., 2023; Li et al., 2015; Bengio et al., 2014; Maiorca et al., 2024; Huh et al., 2024; Guth et al., 2024; Chang et al., 2022; Conneau et al., 2018; Tsitsulin et al., 2020; Nejatbakhsh et al., 2024). This correspondence becomes stronger in wide and large networks (Barannikov et al., 2022; Morcos et al., 2018; Somepalli et al., 2022). Leveraging these aligned embeddings, a simple linear transformation often suffices to map one network’s latent space onto another’s, enabling techniques such as model stitching, where components from different networks can be interchanged with minimal loss in performance (Fumero et al., 2024; Bansal et al., 2021; Csiszárik et al., 2021). In practice, aligning two independently learned latent spaces often requires only a linear transformation, which achieves comparable downstream task performance (Moayeri et al., 2023; Merullo et al., 2023; Maiorca et al., 2024; Lähner and Moeller, 2024).
Latent space geometry.   Early work on the geometry of deep latent representations focused on autoencoders, where the decoder’s mapping from latent to data space induces a natural pullback metric under the assumption that the ambient space is Euclidean (Shao et al., 2018; Tosi et al., 2014; Arvanitidis et al., 2018). The Riemannian viewpoint allows one to compute geodesic paths and meaningful distances that respect the manifold structure of the learned embedding. Subsequent research has introduced computationally efficient approximations, such as energy-based proxies, and extended these ideas to estimate local curvature for improved interpolation and sampling (Chen et al., 2019; Chadebec and Allassonnière, 2022; Loaiza-Ganem et al., 2024; Arvanitidis et al., 2021, 2022a). In the context of discriminative models, one can obtain a Riemannian metric primarily using two approaches (Grosse, 2022), either by pulling back the Fisher Information Matrix (Amari, 2016; Arvanitidis et al., 2022b) or by assuming a Euclidean geometry on the output space and pulling back the $L2$ metric. Interestingly, one can obtain some identifiability guarantees by taking geometry into consideration (Syrota et al., 2025).
@@ 3.1 Notation and background
Neural networks (NNs) are parametric functions $F_{\theta}$, composed of an encoding map and a decoding map, represented as $F_{\theta}=D_{\theta_{2}}\circ E_{\theta_{1}}$. The encoder $E_{\theta_{1}}:\mathcal{X}\mapsto\mathcal{Z}$ generates a latent representation $\bm{z}=E_{\theta_{1}}(\bm{x})$, where $\bm{x}\in\mathcal{X}$ is mapped from the input domain $\mathcal{X}$ to the latent space $\mathcal{Z}$. The decoder $D_{\theta_{2}}$ is responsible for performing the task at hand, such as reconstruction or classification. For simplicity, we omit the parameter dependence ($\theta$) in our notation moving forward. For any single module $E$ (or equivalently $D$), we use $E_{\mathcal{X}}$ to denote that the module $E$ was trained on the domain $\mathcal{X}$. In the next sections, we will provide the necessary background to introduce our method.
@@
Latent space communication.  Given a pair of domains $(\mathcal{X},\mathcal{X}^{\prime})$, a pair of neural models trained on them $(F_{\mathcal{X}}^{1},F_{\mathcal{X^{\prime}}}^{2})$ and a partial correspondence between the domains $\Gamma:\mathcal{A}_{\mathcal{X}}\mapsto\mathcal{A}_{\mathcal{X^{\prime}}}$ where $\mathcal{A}_{\mathcal{X}}\subset\mathcal{X}$ and $\mathcal{A}_{\mathcal{X^{\prime}}}\subset\mathcal{X}^{\prime}$, the problem of latent space communication is the one of finding a full correspondence $\Lambda:E^{1}(\mathcal{X})\mapsto E^{2}(\mathcal{X}^{\prime})$ between the two domains, from $\Gamma$. In a simplified setting, for example two models trained with different initialization or architectures on the same data, $\mathcal{X}=\mathcal{X}^{\prime}$ and the correspondence is the identity. When $\mathcal{X}\neq\mathcal{X}^{\prime}$ the problem recovers the multimodal setting.
@@
Relative representations.  The relative representations framework (Moschella et al., 2023) provides a straightforward approach to represent each sample in the latent space according to its similarity to a set of fixed training samples, denoted as anchors. Representing samples in the latent space as a function of the anchors corresponds to transitioning from an absolute coordinate frame into a relative one defined by the anchors and the similarity function. Given a domain $\mathcal{X}$, an encoding function $E_{\mathcal{X}}:\mathcal{X}\to\mathcal{Z}$, a set of anchors $\mathcal{A}_{\mathcal{X}}\subset\mathcal{X}$, and a similarity or distance function $d:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$, the relative representation for a sample $\bm{x}\in\mathcal{X}$ is:
$RR(\bm{z};\mathcal{A}_{\mathcal{X}},d)=\bigoplus_{\bm{a}_{i}\in\mathcal{A}_{\mathcal{X}}}d(\bm{z},E_{\mathcal{X}}(\bm{a}_{i})),$
@@
where $\bm{z}=E_{\mathcal{X}}(\bm{x})$, and $\bigoplus$ denotes row-wise concatenation. In the original method (Moschella et al., 2023), $d$ corresponds to cosine similarity. This choice induces a representation invariant to angle-preserving transformations. In this work, our focus is to leverage the intrinsic geometry of latent spaces to employ a metric that captures isometric transformations between data manifolds.
@@
Latent space geometry.  For the latent space of a neural network, it is generally hard to reason about its Riemannian structure. However, it is often easier to assign a Riemannian structure to the output space. As such, one can define a pullback metric from the output space to the latent space, which is a standard operation in Riemannian geometry (see Ch.2.4 of Do Carmo and Flaherty Francis (1992)).
@@
Formally, the decoder $D:\mathcal{Z}\mapsto\mathcal{Y}$ takes as input a latent representation $\bm{z}\in\mathcal{Z}$ and outputs $y$. Given a Riemannian metric defined on $y$ as $G_{\mathcal{Y}}(y)$, one can obtain the Riemannian metric at $\bm{z}$ as:
$G_{\mathcal{Z}}(\bm{z})=\left(\frac{\partial\bm{y}}{\partial\bm{z}}\right)^{\!\!\top}G_{\mathcal{Y}}(\bm{y})\left(\frac{\partial\bm{y}}{\partial\bm{z}}\right)=J_{D}(\bm{z})^{\top}G_{\mathcal{Y}}(\bm{y})J_{D}(\bm{z}),$
@@
where $J_{D}(\bm{z})$ is the Jacobian of $D$ at $\bm{z}$. The metric tensor $G_{\mathcal{Y}}$ is useful to compute quantities such as lengths, angles, and areas on $\mathcal{M}$. Given a smooth curve $\bm{\gamma}:[a,b]\mapsto\mathcal{M}$, its arc length is defined as:
$L(\bm{\gamma})$ $=\int^{b}_{a}\sqrt{\bm{v}(t)^{\top}G_{\mathcal{Y}}(\bm{\gamma}(t))\bm{v}(t)^{\top}}\mathrm{d}t,$ (1)
@@
where $\bm{v}(t)=\dot{\bm{\gamma}}(t)$. A slight variation of the above functional gives the geodesic energy $\mathcal{E}$ of $\bm{\gamma}$ (Arvanitidis et al., 2018; Shao et al., 2018)
$\mathcal{E}(\bm{\gamma})$ $=\frac{1}{2}\int^{b}_{a}\bm{v}(t)^{\top}G_{\mathcal{Y}}(\bm{\gamma}(t))\bm{v}(t)^{\top}\mathrm{d}t.$ (2)
@@
Both can be discretized and approximated in practice using finite difference approaches (Yang et al., 2018; Shao et al., 2018). Geodesics minimize both the length and the energy, where for optimization the latter is usually preferred for numerical stability (Hauberg, 2025). These quantities have the property of being invariant to certain reparametrizations, as formalized in the following proposition:
Proposition 3.1.
@@
Let $\bm{\gamma}:[0,1]\to\mathcal{M}$ be a smooth curve on a Riemannian manifold $(\mathcal{M},G)$, and let $(\mathcal{M}^{\prime},G^{\prime})$ be a reparameterization of the manifold and $\varphi:[0,1]\to[0,1]$ a smooth, strictly increasing reparametrization of $\gamma$. Setting $\bm{\gamma}^{\prime}(\tau)=\bm{\gamma}\bigl(\varphi(\tau)\bigr)$ the Riemannian length and energy of $\bm{\gamma}$ are invariant under reparameterizations of the manifold:
$\mathcal{E}[\bm{\gamma}]$ $=\frac{1}{2}\int_{0}^{1}\bigl\lVert\tfrac{d\bm{\gamma}^{\prime}}{d\tau}\bigr\rVert^{2}_{G}\,d\tau=\frac{1}{2}\int_{0}^{1}\bigl\lVert\tfrac{d\bm{\gamma}^{\prime}}{d\tau}\bigr\rVert^{2}_{G^{\prime}}\,dt,$
$L[\bm{\gamma}]$ $=\int_{0}^{1}\bigl\lVert\tfrac{d\bm{\gamma}^{\prime}}{d\tau}\bigr\rVert_{G}\,d\tau=\int_{0}^{1}\bigl\lVert\tfrac{d\bm{\gamma}^{\prime}}{d\tau}\bigr\rVert_{G^{\prime}}\,dt.$
Furthermore, the Riemannian arc length of $\bm{\gamma}$ is invariant under reparametrizations $\bm{\gamma}^{\prime}$ on $\mathcal{M}$:
$L[\bm{\gamma}^{\prime}]=\int_{0}^{1}\bigl\lVert\tfrac{d\bm{\gamma}^{\prime}}{d\tau}\bigr\rVert_{G}\,d\tau=\int_{0}^{1}\lVert\tfrac{d\bm{\gamma}}{dt}\rVert_{G}\,dt=L[\bm{\gamma}].$
We provide the proof in Appendix A.1.1.
@@ 3.2 Relative geodesics representations
1:Sample $\bm{x}\in\mathcal{X}$, anchors $\mathcal{A}_{\mathcal{X}}$, encoder $E$, decoder $D$, distance $d_{\mathcal{X}}$ induced by metric $G_{\mathcal{Y}}$, steps NN, step size $\Delta t$, mode $\in\{\texttt{energy},\texttt{distance}\}$
[…]
@@
5:  $\bm{z}_{a}\leftarrow E(a)$,  $d\leftarrow 0$
6:  for $j=1$ to NN do
7:   $\bm{\gamma}_{j}\leftarrow(1-\tfrac{j}{N})\bm{z}+\tfrac{j}{N}\bm{z}_{a}$
8:   $\bm{\gamma}_{j-1}\leftarrow(1-\tfrac{j-1}{N})\bm{z}+\tfrac{j-1}{N}\bm{z}_{a}$
9:   $\bm{v}\leftarrow D(\bm{\gamma}_{j})-D(\bm{\gamma}_{j-1})$
10:   $G\leftarrow G_{\mathcal{Y}}(D(\bm{\gamma}_{j}))$
11:   $s\leftarrow\bm{v}^{\top}G\bm{v}$
@@
From a differential geometry perspective, the problem of latent space communication can be interpreted as finding a transformation between the data manifolds $\mathcal{M}_{1},\mathcal{M}_{2}$ approximated by two neural models $F_{\mathcal{X}}^{1},F_{\mathcal{X^{\prime}}}^{2}$. The relative representation framework captures this transformation implicitly if equipped with the right metric: we posit that a natural candidate for this metric is the geodesic distance defined on $\mathcal{M}_{1},\mathcal{M}_{2}$, respectively. This choice makes the relative representations invariant to isometric transformation $T$ of the manifolds $\mathcal{M}_{1},\mathcal{M}_{2}$. However, for high-dimensional problems, the high cost of computing the geodesic (corresponding to minimizing Eq. 2) makes this impractical (Shao et al., 2018; Chen et al., 2019). Furthermore, one can argue against directly using the latent geometry induced by deterministic models from a theoretical perspective (Hauberg, 2019), as it may result in undesirable properties, for example the geodesics going outside of the data manifold.
@@
We therefore approximate the geodesic quantities by directly considering the energy (or the length) of the straight line (in the Euclidean sense) connecting representations in the latent space:
$RR^{geo}(\bm{z};\mathcal{A}_{\mathcal{X}})=\bigoplus_{\bm{a}_{i}\in\mathcal{A}_{\mathcal{X}}}\mathcal{E}(\tilde{\bm{\gamma}}_{\alpha}(\bm{z},E_{\mathcal{X}}(\bm{a}_{i}))),$
where $\tilde{\bm{\gamma}}_{\alpha}(\bm{z}_{1},\bm{z}_{2})=(1-\alpha)\bm{z}_{1}+\alpha\bm{z}_{2}$ is the convex combination between the points $\bm{z}_{1},\bm{z}_{2}$. The approximation gives a natural upper bound to the geodesic distance: for $\bar{\bm{\gamma}}$ it can be shown to relate to the arc length of a curve defined in Eq. 1 and the energy in Eq. 2 using the following bounds:
[…]
The proof is in Appendix A.1.2. Moreover the approximation is far more efficient to compute, without requiring minimization of equation 2, and is accurate, as empirically verified in Figure 2.
Discretization.  When the step size is small enough, energy and arc length in the latent space as in Equations 1,2 can be approximated by by their counterpart on the output space using discretized finite difference schemes (Shao et al., 2018):
$\mathcal{E}(\bm{\gamma})$ $=\sum_{i=1}^{N}E_{i}=\frac{1}{2}\sum_{i=1}^{N}\bm{v}(t_{i})^{\top}G(t_{i})\bm{v}(t_{i})\Delta t,$ (4)
$L(\bm{\gamma})$ $=\sum_{i=1}^{N}d_{i}=\sum_{i=1}^{N}\sqrt{\bm{v}(t_{i})^{\top}G(t_{i})\bm{v}(t_{i})}\Delta t,$ (5)
where $\Delta t=\frac{1}{N}$, with $N$ being the number of discretization steps. For Euclidean geometry, the geodesic arc lengths are given in closed form as the geodesics are straight lines. Unlike the energy, the curve length is invariant under reparametrizations (proposition 3.1). As such we focus on the curve length in our experiments. The resulting algorithm is summarized in Algorithm 1. In practice, with specific choice of $G_{\mathcal{Y}}$ one can avoid approximating the distance between $D(\bm{\gamma}_{j},\bm{\gamma}_{j-1})$ explicitly using $G_{\mathcal{Y}}$ by directly calculating the distance or energy between $\bm{\gamma}_{j}$ and $\bm{\gamma}_{j-1}$ on $\mathcal{Y}$.
@@
Approximate geodesic energies.  Our choice comes with three advantages: (i) efficiency: avoiding minimization of Eq.2 the computation for every sample reduces to a single forward pass for every discretization step $\bm{\gamma}$ and for each anchor, resulting in overall complexity of $O(TA)$ forward passes of the decoder, where $A$ the number of anchors is the number of discretization steps. (ii) Directly using the arc length ensures invariance to reparametrizations of the manifold, matching our assumptions. (iii) As we only need reasonably accurate estimates of the arc lengths rather than the geodesic trajectory, the approach is accurate. Specifically, to assess how close the straight line energy approximation (2) is to the true geodesic energies, we first encoded 100 samples (10 per class, sorted by label) from MNIST (Deng, 2012) and CIFAR-10 using a simple convolutional autoencoder (architecture detailed in Appendix A.3.2). We then computed pairwise geodesic energy matrices over these latent representations using both methods, and the results are displayed in Fig. 2. Visually, both energy matrices exhibit the same block-diagonal structure, mainly due to belonging to the same class, and clustering patterns. Numerically, their Spearman rank correlation exceeds 0.99 with only 8 discretization points (see Appendix A.3.5 for correlation results across different numbers of discretization steps and for implementation details).
[…]